# Memory preservation and cooperative shielding in complex quantum networks

### v2

@@ 2.2 Role of the anisotropy (ZZ-term interaction)
Since the levels are degenerate, we should in principle compute (at first order) the eigenvalues of the $\hat{\mathcal{W}}$ perturbation restricted to the considered eigenspace. However, it can be easily checked by inspection that $\hat{\mathcal{W}}$ is already expressed in a diagonal form, due to the definitions of the computational basis $\mathcal{P}$. Therefore, we only need a restricted number of matrix elements. For any form of the $\Delta_{ij}$ coefficients one can cast the action of the perturbation operator on a given state in the computational basis $|\boldsymbol{\sigma}_{q}\rangle$, with $\boldsymbol{\sigma}_{q}$ expressing the spin polarizations vector in each site, the index $q$ denoting the specific Pauli string. We thus obtain:
$\langle\boldsymbol{\sigma}_{q}|\hat{\mathcal{W}}|\boldsymbol{\sigma}_{q}\rangle=\sum_{i<j}\Delta_{ij}\frac{\text{sgn}(i,q)\text{sgn}(j,q)}{4}\,$, (9)
@@
We can note how the shifts in energy depend on the pairwise products of the spins’ directions in the basis vector. In order to simplify our analytical computation, we consider now nearest-neighbor interaction, in other words $\Delta_{ij}=\Delta_{0}$ if dist(i,j)=1(i,j)=1 and vanishing otherwise. Under this condition, the energy shifts read:
$\langle\boldsymbol{\sigma}_{q}|\hat{\mathcal{W}}|\boldsymbol{\sigma}_{q}\rangle=\frac{\Delta_{0}}{4}\sum_{i=1}^{L}\text{sgn}(i,q)\text{sgn}(i+1,q)\quad,\quad L+1\equiv 1\,\text{(PBC)}\,.$ (10)
@@
For more general interactions, going back to Eq. (9), we consider the case of the first two neighbors having non-negligible interactions and with PBC:
$\langle\boldsymbol{\sigma}_{q}|\hat{\mathcal{W}}|\boldsymbol{\sigma}_{q}\rangle=\frac{\Delta_{1}}{4}\sum_{i=1}^{L}\text{sgn}(i,q)\text{sgn}(i+1,q)+\frac{\Delta_{2}}{4}\sum_{i=1}^{L}\text{sgn}(i,q)\text{sgn}(i+2,q)\,$,
@@
Without loss of generality, we can hence turn to examine in detail the case of nearest-neighbor interaction for the perturbation. The shift reflects the fact that if there is a total number of $L$ couples (with PBC) and $b$ is the number of pairs with opposite spin directions in a given Pauli string, the sum (10) corresponds to the number of consecutive pairs with spins pointing in the same direction after subtracting those with opposite spin, that amounts to (Lb)b=L2b(L-b)-b=L-2b. By way of example, let us compare the resulting effect for the cases of one and two excitations. In the former case, there are only two out of the $L$ couples of spins that have opposite directions. Thus, the summation in (10) amounts to L4L-4 for each vector state of the restricted computational basis $\mathcal{P}_{1}$. This means that the shift is the same for both levels: importantly the perturbation has no overall effect on the local $z$-magnetization frequency. In the latter case, one can distinguish two possibilities for the Pauli strings, as depicted in table 1. According to the considered vector, one can get different values for the summation in eq. (10) and hence for the energy shift.
@@
We thus see that after introducing a given number of excitations, the maximum shift interval in (11) is proportional to the number of excitations in the quantum spin network. As a result, the quantum network develops an uncertainty in the encoding of the number of excitations, that scales linearly with the number of excitations itself. The physical understanding of this network property is that the anisotropic interaction between nearest-neighbor spins shifts the original energy levels, thus creating enlarged spectrum from degeneracy breaking with further and wider peaks.
This result opens the door to applications in quantum technologies exploiting the memory-preserving property of the XXZ quantum network by engineering the ZZ anisotropy to obtain a constant relative uncertainty in the encoding of the number of excitations.
@@
Interestingly, this result can also be linked to the description of information processing properties of the out-of-equilibrium XXZ network with all-to-all connectivity analyzed in [5, 14]. There, an ideal-observer analysis of the power spectra had shown that the uncertainty with which the quantum network encodes the number of frequency peaks - and therefore the number of excitations - linearly scales with the number itself. In neuroscience, this is the content of Weber’s law describing the perception of the number of items in a space, or numerosity perception. This phenomenology is experimentally observed under very general conditions, no matter the perception channel (i.e. visual or auditory), nor the way the items are presented [72]. Importantly, neuroscience experiments indicate that the perception of number is interconnected to the perception of time and space intervals [77].
The overall bottom line is that even in the XXZ quantum network the emerging dynamical frequencies - or else the numerosity counting itself- comes from the number of consecutive gaps between the original levels. In fact, for small enough perturbations, anisotropy is found to just inject an uncertainty due to degeneracy lifting and the consequent spreading of the bands, the latter depending itself on the number of excitations introduced.
@@
With this physical understanding at hand, we can now turn our attention to the transport properties of more general quantum networks, where we release either the constraint of constant coupling intensity of the links in the regular network, or the regular architecture itself by considering complex quantum networks. This is the subject of the next Section 3, where we explore the extent to which the space and time memory-preserving property of the regular XXZ network excitations can be engineered in more general networks; We will focus, in particular, to the case of one initial excitation or, more generally, in the m=L/2+1m=-L/2+1 total $z$-magnetization sector.
3 Transport properties in generalized quantum networks
@@
So far we built a theoretical understanding about the out-of-equilibrium dynamical behavior of the closed quantum spin network governed by the XXZ Hamiltonian, for a network with all-to-all connectivity architecture of equally weighted links. For this case, we could resort to semi-analytical means. In this section, we start from the understanding developed in Section 2 to numerically investigate the role played by (i) the interaction graph in a complex quantum system architecture, and (ii) the range of the interactions according to a power law when all sites are connected between them. For simplicity, we stick to the case of one single excitation, having developed in Section 2 an intuition for multiple excitations: we will come back to this point in the concluding Section 4. As an example for complex quantum network, we choose the small-world architecture: besides being per se relevant for diverse problems, we are here interested to its property of well interpolating between the cases of regular and random graphs after tuning only two parameters.
@@
We analyze the underlying physics by specializing our analysis to a localization quantifier called Inverse Participation Ratio (IPR), see Appendix B.
3.1 Small-world networks
@@
A number of real-life networks relevant to diverse phenomena, ranging from biology to social sciences, are characterized by a high level of clustering and, at the same time, by a low average distance between the vertices: this is the so-called “small-world phenomenon” that illustrates e.g. social networks and the World Wide Web [78].
@@
Such a network can be built from a regular-lattice graph where each node is linked to its first $k$ neighbors, and then introducing a rewiring probability $\beta$ that one edge connects the starting node to a different one. The parameter $2k$ drives the level of network complexity, as it denotes the number of initial connections for each site. The parameter $\beta$ governs the level of graph’s randomness, bridging between the regular-lattice case with $\beta=0$ and the fully random graph with $\beta=1$. The all-to-all connected graph is obtained in the limit $\beta=0$ and 2k/(L1)=12k/(L-1)=1. The construction mechanism is illustrated in Fig. 4(a)(a).
@@
We complete this introductory section by recalling the main qualifiers and quantifiers for the networks architectures considered in complex-network theory, that will be used in the following. For details, we refer to the Appendix C. One first connectivity quantifier is the node degree, denoted by the word “deg”, that is the number of edges linking one node to others. The coupling range is quantified by the topological distance distij, counting the number of links connecting two sites in their shortest path: a global distance quantifier $\mathcal{D}$ can then be usefully introduced, as the average of the topological distance over all the pairs of nodes. Finally, the tendency to aggregate is quantified by the global clustering coefficient $\mathcal{C}$, that is the average proportion of total triangles (meaning closed loops of three vertices) within a given neighborhood, as related to the maximum number of allowed connections according to the local degree. All-to-all connected graphs are the limiting case of low topological distance, i.e. tending to 1, and high clustering coefficient, i.e. tending to 1; in the opposite limiting case of nearest-neighbor networks both local and global clustering are 0, since no triangles between nodes can be observed, while average distance linearly scales as the number of sites $L$.
3.2 Memory effects in small-world networks
@@
Here we analyze the memory effect discussed in Section 2.1 in small-world networks, starting with a single excitation, i.e. a single spin oriented along $z$ while all the other spins are along the opposite direction. We use the same technique of Section 2.1 in which we take the Fourier transform of $S^{z}_{\text{initial}}(t)$. In presence of a single excitation only one high-energy resonance (whose energy scale with the system size) is present in the dynamics for all-to-all interaction, allowing us to count the number of excitations. Here we analyze how the Fourier spectrum changes as we decrease the connectivity. Interestingly, even if the number resonances increases as we decrease the connectivity, the high-energy resonances persists even for lower connectivity, see Fig. 4(c)(c). This result could allow to extend the notion of memory effect to generic small-word network. Indeed, it shows that signatures of the memory-preserving effect survive even for less connected network. Clearly, it would be interesting to study if the memory-preserving effect persists in small-world networks also in the many excitation case. This will be the topic of a future work.
@@
Figure 4: Construction of small-world graphs and their memorization properties. $(a)$ Small-world network construction (see text). Top to bottom: a regular graph with $L=7$ sites and $k=2$ nearest-neighbor connections, and different rewired links (green and orange) with respect to the original graph, according to the value of the probability $\beta$. (b)(e)(b)-(e) Examples of intensity spectra for average $z$-magnetization at the initial site for one quenched excitation along the $z$-axis in four different networks with $L=33$ sites. $(b)$ All-to-all network, with delta-like peaks at frequencies $\omega^{*}=\pm JL/2$=33, see Eq. 4. $(c)$ Small-world, highly-connected graph with $(k,\beta)$=(15,0.5), possessing additional low-frequency components. $(d)$ Small-world, highly-connected graph with $(k,\beta)$=(13,0.5), having small high-frequency contributions. $(e)$ Sparsely connected small-world with $(k,\beta)$=(1,0.2), having no high-frequency peaks. The data has been detrended before the Fourier transform by removing the zero-frequency component.
3.3 Inverse Participation Ratio
@@
defined as the sum of the squared trace elements of the density matrix $\hat{\mathcal{\varrho}}$ over single-excitation localized states $\{|s\rangle\}=\{|\downarrow\downarrow\dots\downarrow\uparrow\downarrow\dots\downarrow\rangle\}$, where the $\uparrow$ in the previous Pauli string is a spin up (along $z$) located at site $s$; this basis has dimension $L$ and constitutes one of the subspaces of the total $z$-magnetization operator $\hat{\mathcal{M}}_{z}$ with which $\hat{\mathcal{H}}$ can be block-diagonalized. While we refer to Appendix B for a complete definition, we here highlight that by definition the IPR values range from $1/L$ for the complete delocalized state, i.e. infinite-temperature case, to $1$ for the perfect localization on a single eigenstate. If needed, generalizations of this quantifiers to the case of multiple excitations can be introduced, as in [4].
@@
Before proceeding, it is useful to briefly discuss the form of the IPR function for a single excitation in the all-to-all case, as in Section 2: following the theoretical discussion in Section 2.1, one can easily derive the analytical function as the sum of a constant term and two cosine functions, oscillating with frequencies $\omega^{*}$ and $2\omega^{*}$, as $\omega^{*}$ is the same frequency of the local $z$-magnetization signals that is the energy gap between the levels $|l=L/2,m=-l/2+1\rangle$ and $|l=L/2-1,m=-L/2+1\rangle$, see Eq. 4. Therefore, in the following we will also show the IPR average for the all-to-all case over a temporal window $2\pi/\omega^{*}$. The way delocalization happens in small-world networks is strongly dependent on the initial conditions, which for our case translates to the site of the first excitation and its degree; this is the focus of Section 3.4.
3.4 Dependence on initial conditions

### v3

@@ II.2 Role of the anisotropy (ZZ-term interaction)
Since the levels are degenerate, we should in principle compute (at first order) the eigenvalues of the $\hat{\mathcal{W}}$ perturbation restricted to the considered eigenspace. However, it can be easily checked by inspection that $\hat{\mathcal{W}}$ is already expressed in a diagonal form, due to the definitions of the computational basis $\mathcal{P}$. Therefore, we only need a restricted number of matrix elements. For any form of the $\Delta_{ij}$ coefficients one can cast the action of the perturbation operator on a given state in the computational basis $|\bm{\sigma}_{q}\rangle$, with $\bm{\sigma}_{q}$ expressing the spin polarizations vector in each site, the index $q$ denoting the specific Pauli string. We thus obtain:
$\langle\bm{\sigma}_{q}|\hat{\mathcal{W}}|\bm{\sigma}_{q}\rangle=\sum_{i<j}\Delta_{ij}\frac{\text{sgn}(i,q)\text{sgn}(j,q)}{4}\,$, (9)
@@
We can note how the shifts in energy depend on the pairwise products of the spins’ directions in the basis vector. In order to simplify our analytical computation, we consider regular networks where each node is connected to its first $k$ neighbors; arbitrary topologies introduced by the perturbation can be numerically computed from Eq. (9). Starting with first, nearest-neighbors case, i.e. $\Delta_{ij}=\Delta_{0}$ if dist(i,j)=1(i,j)=1 and vanishing otherwise, the energy shifts read:
$\langle\bm{\sigma}_{q}|\hat{\mathcal{W}}|\bm{\sigma}_{q}\rangle$ $=\frac{\Delta_{0}}{4}\sum_{i=1}^{L}\text{sgn}(i,q)\text{sgn}(i+1,q),$ (10)
$L+1$ $\equiv 1\,\text{(PBC)}.$
@@
Instead, in the case of the first two neighbors having non-negligible interactions and within PBC:
$\langle\bm{\sigma}_{q}|\hat{\mathcal{W}}|\bm{\sigma}_{q}\rangle$ $=\frac{\Delta_{1}}{4}\sum_{i=1}^{L}\text{sgn}(i,q)\text{sgn}(i+1,q)$
$\quad+\frac{\Delta_{2}}{4}\sum_{i=1}^{L}\text{sgn}(i,q)\text{sgn}(i+2,q),$ (11)
@@
The energy shift formula (11) being similar in form to (10), we can hence turn to examine in detail the case of nearest-neighbor interaction for the perturbation. The shift reflects the fact that if there is a total number of $L$ couples (with PBC) and $b$ is the number of pairs with opposite spin directions in a given Pauli string, the sum of Eq.(10) corresponds to the number of consecutive pairs with spins pointing in the same direction after subtracting those with opposite spin, that amounts to (Lb)b=L2b(L-b)-b=L-2b. By way of example, let us compare the resulting effect for the cases of one and two excitations. In the former case, there are only two out of the $L$ couples of spins that have opposite directions. Thus, the summation in Eq.(10) amounts to L4L-4 for each vector state of the restricted computational basis $\mathcal{P}_{1}$. This means that the shift is the same for both levels: importantly the perturbation has no overall effect on the local $z$-magnetization frequency. In the latter case, one can distinguish two possibilities for the Pauli strings, as depicted in table 1. According to the considered vector, one can get different values for the summation in Eq.(10) and hence for the energy shift.
@@
We thus see that after introducing a given number of excitations, the maximum shift interval in 12 is proportional to the number of excitations in the quantum spin network. As a result, the quantum network develops an uncertainty in the encoding of the number of excitations, that scales linearly with the number of excitations itself. The physical understanding of this network property is that the anisotropic interaction between nearest-neighbor spins shifts the original energy levels, thus creating enlarged spectrum from degeneracy breaking with further and wider peaks.
This result can provide useful insights in the design of quantum technology applications embodying multiplicative noise, i.e. proportional to signal, thus leading to equal information per logarithmic interval of the variable [gardiner2004stochastic, nielsen2000quantum]. This can be relevant for the design of self-similar quantum networks or of non-classical states of matter for quantum metrology. Here, fluctuations due to phase diffusion, such as with collective spin states, cause the nature of noise to be multiplicative and represent the realistic limit to ideal quantum-sensitivity enhancement [demkowicz-dobrzanski2012metrology].
@@
Interestingly, this result can also be linked to the description of information processing properties of classical complex systems, performed with some quantum advantage by means of quantum network models. One such example concerns the numerosity perception in neuroscience, that is the ability of humans and mammals of counting the number of items in a given space without cognitively counting, performing with an about 15% relative uncertainty in a quite large range of items number, known as Weber’s law [dehaene2003weberlaw]. This phenomenology is experimentally observed under very general conditions, no matter the perception channel (i.e. visual or auditory), nor the way the items are presented [burr2008visual]. Importantly, neuroscience experiments indicate that the perception of number is interconnected to the perception of time and space intervals [Morrone:2007], challenging neuroscience understanding. In [YagoMalo2023, yagomalo2024review], the mapping of the problem into an out-of-equilibrium XXZ network with all-to-all connectivity has been shown to account for the Weber’s law in effective and efficient manner, whereas conventional artificial neural networks poorly perform [nasr2019deepneural]. There, an ideal-observer analysis of the power spectra had shown that the uncertainty with which the quantum network encodes the number of frequency peaks - and therefore the number of excitations - linearly scales with the number itself. The analytical understanding that we have built up in the present work provides neuroscientists with useful insights and can be promising to link numerosity perception with time and space-intervals perception.
@@
With this physical understanding at hand, we can now turn our attention to the transport properties of more general quantum networks, where we release either the constraint of constant coupling intensity of the links in the regular network, or the regular architecture itself by considering complex quantum networks. This is the subject of the next Section III, where we explore the extent to which the space and time memory-preserving property of the regular XXZ network excitations can be engineered in more general networks; We will focus, in particular, to the case of one initial excitation or, more generally, in the m=L/2+1m=-L/2+1 total $z$-magnetization sector.
III Transport properties in generalized quantum networks
@@
So far we built a theoretical understanding about the out-of-equilibrium dynamical behavior of the closed quantum spin network governed by the XXZ Hamiltonian, for a network with all-to-all connectivity architecture of equally weighted links. For this case, we could resort to semi-analytical means. In this Section, we start from the understanding developed in Section II to numerically investigate the role played by (i) the interaction graph in a complex quantum system architecture, and (ii) the range of the interactions according to a power law when all sites are connected between them. For simplicity, we stick to the case of one single excitation, having developed in Section II an intuition for multiple excitations: we will come back to this point in the concluding Section IV. As an example for complex quantum network, we choose the small-world architecture: besides being per se relevant for diverse problems, we are here interested to its property of well interpolating between the cases of regular and random graphs after tuning only two parameters.
@@
We analyze the underlying physics by specializing our analysis to a localization quantifier called Inverse Participation Ratio (IPR), see Appendix C. The analysis that follows will follows on the XX part of Hamiltonian (1), i.e. all coefficients $\Delta_{ij}$ will be set to 0.
III.1 Small-world networks
@@
A number of real-life networks relevant to diverse phenomena, ranging from biology to social sciences, are characterized by a high level of clustering and, at the same time, by a low average distance between the vertices: this is the so-called “small-world phenomenon” that illustrates e.g. social networks and the World Wide Web [watts1998smallworld].
@@
Such a network can be built from a regular-lattice graph where each node is linked to its first $k$ neighbors; going clockwise through all nodes, for all $k$ links situated on the left of it we rewire it with probability $\beta$ to a different one. For the types of network we are considering, $k$ can vary from 1 to $\lfloor L/2\rfloor$; related to it, the parameter $\bar{g}=2k$, ranging from 2 to L1L-1 by steps of 2, drives the level of network complexity, as it denotes the number of initial connections for each site and also the average degree defined below, due to the way these networks are constructed. The parameter $\beta$ governs the level of graph’s randomness, bridging between the regular-lattice case with $\beta=0$ and the fully random graph with $\beta=1$. The all-to-all connected graph is obtained in the limit $\bar{g}=L-1$, while the nearest-neighbor case has $\bar{g}=2$. The construction mechanism is illustrated in Fig. 4(a)(a). Throughout the rest of the article we treat the case of odd $L$ for ease of notation without loss of generality.
@@
We complete this introductory Section by recalling the main qualifiers and quantifiers for the networks architectures considered in complex-network theory, that will be used in the following. For details, we refer to the Appendix D. One first connectivity quantifier is the node degree, denoted by the word “deg”, that is the number of edges linking one node to others. The coupling range is quantified by the topological distance distij, counting the number of links connecting two sites in their shortest path: a global distance quantifier $\mathcal{D}$ can then be usefully introduced, as the average of the topological distance over all the pairs of nodes. Finally, the tendency to aggregate is quantified by the global clustering coefficient $\mathcal{C}$, that is the average proportion of total triangles (meaning closed loops of three vertices) within a given neighborhood, as related to the maximum number of allowed connections according to the local degree. All-to-all connected graphs are the limiting case of low topological distance, i.e. tending to 1, and high clustering coefficient, i.e. tending to 1; in the opposite limiting case of nearest-neighbor networks both local and global clustering are 0, since no triangles between nodes can be observed, while average distance linearly scales as the number of sites $L$.
III.2 Memory effects in small-world networks
@@
Here we analyze the memory effect discussed in Section II.1 in small-world networks, starting with a single excitation, i.e. a single spin oriented along $z$ while all the other spins are along the opposite direction. We use the same technique of Section II.1 in which we take the Fourier transform of $S^{z}_{\text{initial}}(t)$. In presence of a single excitation only one high-energy resonance (whose energy scale with the system size) is present in the dynamics for all-to-all interaction, allowing us to count the number of excitations. Here we analyze how the Fourier spectrum changes as we decrease the connectivity. Interestingly, even if the number resonances increases as we decrease the connectivity, the high-energy resonances persists even for lower connectivity, see Fig. 4(c)(c). This result could allow to extend the notion of memory effect to generic small-word network. Indeed, it shows that signatures of the memory-preserving effect survive even for less connected network. Clearly, it would be interesting to study if the memory-preserving effect persists in small-world networks also in the many excitation case. This will be the topic of a future work.
@@
Figure 4: Construction of small-world graphs and their memorization properties. $(a)$ Small-world network construction (see text). Top to bottom: a regular graph with $L=7$ sites and $k=2$ nearest-neighbor connections, and different rewired links (green and orange) with respect to the original graph, according to the value of the probability $\beta$. (b)(e)(b)-(e) Examples of intensity spectra for average $z$-magnetization at the initial site for one quenched excitation along the $z$-axis in four different networks with $L=33$ sites. $(b)$ All-to-all network, with delta-like peaks at frequencies $\omega^{*}=\pm JL/2$=33, see Eq.(4). $(c)$ Small-world, highly-connected graph with $(k,\beta)$=(15,0.5), possessing additional low-frequency components. $(d)$ Small-world, highly-connected graph with $(k,\beta)$=(13,0.5), having small high-frequency contributions. $(e)$ Sparsely connected small-world with $(k,\beta)$=(1,0.2), having no high-frequency peaks. The data has been detrended before the Fourier transform by removing the zero-frequency component.
III.3 Inverse Participation Ratio
@@
defined as the sum of the squared trace elements of the density matrix $\hat{\mathcal{\varrho}}$ over single-excitation localized states $\{|s\rangle\}=\{|\downarrow\downarrow\dots\downarrow\uparrow\downarrow\dots\downarrow\rangle\}$, where the $\uparrow$ in the previous Pauli string is a spin up (along $z$) located at site $s$; this basis has dimension $L$ and constitutes one of the subspaces of the total $z$-magnetization operator $\hat{\mathcal{M}}_{z}$ with which $\hat{\mathcal{H}}$ can be block-diagonalized. While we refer to Appendix C for a complete definition, we here highlight that by definition the IPR values range from $1/L$ for the complete delocalized state, i.e. infinite-temperature case, to $1$ for the perfect localization on a single eigenstate. If needed, generalizations of this quantifiers to the case of multiple excitations can be introduced, as in [Civolani2024].
@@
Before proceeding, it is useful to briefly discuss the form of the IPR function for a single excitation in the all-to-all case, as in Section II: following the theoretical discussion in Section II.1, one can easily derive the analytical function as the sum of a constant term and two cosine functions, oscillating with frequencies $\omega^{*}$ and $2\omega^{*}$, as $\omega^{*}$ is the same frequency of the local $z$-magnetization signals that is the energy gap between the levels $|l=L/2,m=-l/2+1\rangle$ and $|l=L/2-1,m=-L/2+1\rangle$, see Eq.(4). Therefore, in the following we will also show the IPR average for the all-to-all case over a temporal window $2\pi/\omega^{*}$. The way delocalization happens in small-world networks is strongly dependent on the initial conditions, which for our case translates to the site of the first excitation and its degree; this is the focus of Section III.4.
III.4 Dependence on initial conditions